Problem 3.062 Your answer is partially correct. Try again. Determine the mean and variance of the random variable with the following probability mass function. Hint: Use the formula for the sum of the infinite series Ra-Taa, where lak1 Give exact answers in form of fraction. 1/5 Variance - 4/15
Answer #1

The given Probability mass function is that of a geometric distribution with probability of success, p = 5/6

Mean of geometric distribution is given by : (1-p) / p

Variance of geometric distribution is given by : (1-p) / p2

Putting the value of p = 5/6 in the above formula, we get :

Mean = 1/5

Variance = 6/25